OpenStudy (anonymous):

Solve the following system of equations. 2a – 3b + c = 10 2a – 2b – 2c = 2 a + 3b + 2c = –1 (2, 3, 4) (1, –2, 2) (3, –1, 0) (–1, 2, 3)

OpenStudy (anonymous):

Can someone please help me?

OpenStudy (anonymous):

is it 1, -2, 2

OpenStudy (anonymous):

ok - add equn 1 to equn. 3 to eliminate b 3a + 3c = 9...............(1) multiply second equn by 1.5: 3a - 3b - 3c = 3 add this to equn. 3: 4a - c = 2..........(2) multiply (2) by 3: 12a - 3c = 6........(3) add (1) and (3): 15a = 15 a = 1 and 4 - c = 2 so c = 2 sub a=1 c=2 in first equn.: 2*1- 3b + 2= 10 3b = 4 - 10 = -6 b = -2 correct option is (1,-2, 2)